JAMES HAS x CHILDREN WITH HIS FIRST WIFE. MARY HAS x+1 CHILDREN WITH HER FIRST HUSBAND .THEY MARY EACH OTHER AND HAVE CHILDREN OF THEIR OWN .THERE ARE TOTAL 24 CHILDRENS. ASSUMING THAT NO TWO CHILDREN  OF SAME PARENTS FIGHT WITH EACH  OTHER. HOW MANY TOTAL FIGHTS THERE WILL BE............. 

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear rahul

let james and mary have y children

x+(x+1) +y=24

 y= 23-2x

so total number of fight =24C2

these fights also include firght among the children of same parents.

so children of james and his first wife can fight =xC2  ways

and children of Mary and her first husband  can fight =x+1C2  ways

 children of james and mary can fight =yC2 = (23-2x)C2 ways

so total number of desired fight

      24C2 -xC2 -x+1C2  -(23-2x)C2 

simplify  -3x2 +45x+23

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin